# A faster FPRAS for #NFA

### 20 Dec 2023 v1

@@ 4. Correctness of Algorithm 2
$=$ $\frac{\textsf{sz}_{i_{0}}}{\sum_{i}^{k}\textsf{sz}_{i}}\times\frac{1}{|T_{i_{0}}|}$
The last equality is because we have assumed $\forall i$, $|S_{i}|\geq t$, so the if condition in line 2 is always satisfied. Hence once $i_{0}$ is sampled in line 2 probability that $\sigma_{0}$ is obtained is exactly $1/|T_{i_{0}}|$, since the set $S_{i_{0}}$ contains elements sampled uniformly (with replacement) from $T_{i_{0}}$. Therefore,
$\frac{1}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\Pr[(\sigma_{0},i_{0})\mbox{ is sampled }]\leq(1+\epsilon_{\textsf{sz}})\frac{1}{\sum_{i=1}^{k}\textsf{sz}_{i}}.$
Let $Y_{j}$ be the random variable denoting whether the counter $Y$ is increased in the $j$th iteration of the loop. Note that $Y_{j}=1$ if the pair $(\sigma_{0},i_{0})$ sampled in the $j$th iteration is in $\mathcal{U}_{1}$. This is because in line 2 it is checked if the sampled pair is in $\mathcal{U}_{1}$.
So, $\mathbb{E}[Y_{j}]=\sum_{(\sigma,i)\in\mathcal{U}_{1}}\Pr[(\sigma,i)\mbox{ is sampled}]$. Thus,
$\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\leq\mathbb{E}[Y_{j}]\leq(1+\epsilon_{\textsf{sz}})\frac{|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$
Note that $Y=\sum_{i}^{t}Y_{i}$. Thus at the end of the algorithm the expected value of $Y$ is between $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}$ and $(1+\epsilon_{\textsf{sz}})^{2}\frac{t|\mathcal{U}_{1}|}{\sum_{i=1}^{k}\textsf{sz}_{i}}$.
Now by Chernoff bound we have
$\Pr[Y\not\in(1\pm\frac{\epsilon}{2})\mathbb{E}[Y]]\leq 2exp(-\frac{\epsilon^{2}\mathbb{E}[Y]}{12}).$
We also know that,
$\mathbb{E}[Y]\geq$ $\frac{t|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}|T_{i}|}{(1+\epsilon_{\textsf{sz}})\sum_{i=1}^{k}\textsf{sz}_{i}}\geq\frac{t\max_{i}\textsf{sz}_{i}}{(1+\epsilon_{\textsf{sz}})^{2}\sum_{i=1}^{k}\textsf{sz}_{i}}$
$\geq$ $12\log(\frac{4}{\delta})$
The last inequality is due to the setting of the parameter $t$. Thus, by Chernoff bound, probability that the output of the algorithm is not between $(1-\frac{\epsilon}{2})\frac{|\mathcal{U}_{1}|}{(1+\epsilon_{\textsf{sz}})}$ and $(1+\frac{\epsilon}{2})(1+\epsilon_{\textsf{sz}})|\mathcal{U}_{1}|$ is at most $\delta/2$. Since $(1-\frac{\epsilon}{2})\geq\frac{1}{1+\epsilon}$, so we have $\Pr[\mathsf{Fail}\ \cap\ \overline{\mathsf{Bad}}]\leq\frac{\delta}{2}$.
Part 2: If for all $i$, $|S_{i}|\geq\texttt{thresh}$
Note that, the else clause in line 2 to 2 is ever reached, if for some $i$, the number of times $i$ is sampled in the $t$ iterations is more than thresh.
@@
Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in line 2) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (lines 2-2). And let the event $\mathsf{Bad}$ be $\cup_{i=1}^{k}\mathsf{Bad}_{i}$. We will now upper bound, $\Pr[\mathsf{Bad}_{i}]$. Let $W_{i}$ be the random variable counting the number times $i$ is sampled. $\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$. By simple application of Chernoff Bound we see that
$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta})\leq exp(-\frac{\Delta\mathbb{E}[W_{i}]}{2}),$
@@
the last inequality is for $\Delta\geq 1$. For our purpose if we put $\Delta=\log(\frac{2k}{\delta})\frac{1}{\mathbb{E}[W_{i}]}$, since $|S_{i}|\geq\texttt{thresh}\geq\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ + $\log(\frac{2k}{\delta})$, we have $\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$. So by union bound $\Pr[\mathsf{Bad}]\leq\delta/2$.
5. Correctness of Algorithm 3
In the following three sections, we prove the three parts of Theorem 2 individually.
5.1. Proof of Theorem 2(1)
@@
Consider an execution of Algorithm 3 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 3. Let $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}$ be the value of the variable $\mathsf{pr_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}$ at level $i$ and let $v$ be the value of $\varphi$ before the function returns. Then we have,
@@
Thus,
(2) $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
Recall that $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus,
$|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
@@
Now, observe that $|\mathcal{L}(P^{0})|=1$ and $P^{\ell}=\{{q}^{\ell}\}$. Thus, under the event $\big{(}\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\big{)}=\texttt{Accurate}\texttt{AppUnion}$, we have
@@
It then follows that
${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=0$
Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}$, we then have
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]\\ =v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell}\cdot p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},\ell-1}\cdots p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},1}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}\end{array}$
5.2. Proof of Theorem 2(2)
@@
$=1{-}\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\gamma_{0}=1{-}\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}$
$\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
5.3. Proof of Theorem 2(3)
@@
    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|$
@@
    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|$.
@@
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\\ \geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
6. Correctness of Algorithm 4
We first establish some preliminaries relevant to our proof.
6.1. Total Variation Distance
The total variation distance is a statistical distance between two discrete valued random variables $X,Y$ over domain $\Omega$ and is defined as follows:
$\mathsf{TV}(X,Y)$ $=$ $\frac{1}{2}\sum_{v\in\Omega}\left|\Pr[X=v]-\Pr[Y=v]\right|$
$=$ $\sum_{v\,:\,\Pr[X=v]\geq\Pr[Y=v]}\left(\Pr[X=v]-\Pr[Y=v]\right)\;.$
We will use the following folklore result (c.f. Lemma 6.3 of  (Vadhan, 2012)). Here, for a random variable $X$ over the sample space $\Omega$, we use the notation ${\mathsf{Pr}}\!\left[E:X\right]$ to denote the probability of the event $E\subseteq\Omega$.
Lemma 6.0.
Let $X$ and $Y$ be the random variables taking values over a measurable set $\Omega$. Let $f:\Omega\to\hat{\Omega}$ be a (possibly randomized) function with measurable co-domain $\hat{\Omega}$. Let $E\subseteq\hat{\Omega}$ be an arbitrary event. Then, the following holds:
$|{\mathsf{Pr}}\!\left[E:{f(X)}\right]-{\mathsf{Pr}}\!\left[E:f(Y)\right]|\leq\mathsf{TV}(X,Y)$
6.2. Proof of Theorem 3
We prove the main result (Theorem 3) by induction on the level $\ell$. For each level $\ell$, we first characterize the accuracy of the computed estimates $\{\mathsf{N}({q}^{\ell})\}_{q\in Q}$ and the quality of the sampled sets $\{\mathsf{S}({q}^{\ell})\}_{q\in Q}$, assuming that these were computed when the samples at the lower levels $11,2,\ldots,\ell-$ 1 are perfectly uniform and have size thresh. After this, we characterize the real computed estimates and samples by arguing that the distance of the corresponding random variables, from those computed using perfectly uniform samples, is small. We will be using the total variation distance and Lemma 1 to structure the proof. We first establish some helpful auxiliary results (Lemma 2 and Lemma 3) towards our main proof. As before, we will often invoke the the probability space obtained by replacing the sets $\mathsf{S}({q}^{j})$ by $\mathsf{U}({q}^{j})$ for all $q\in Q$ and for all $1\leq j\leq i$. In addition to the previously introduced notation ${\mathsf{Pr}}\!\left[;{{\mathsf{U}}_{\leq i}}\right]$, we will also use the notation ${\mathsf{TV}}(X,Y\;;{\mathsf{U}}_{\leq i})$ (resp. ${\mathsf{TV}}(X,Y\;\;\big{|}\;\;\mathcal{F}\;;{\mathsf{U}}_{\leq i})$) to denote the total variational distance between the distribution induced by random variables $X,Y$ (resp. distribution induced by $X,Y$ conditioned on the event $\mathcal{F}$) when the $\mathsf{S}({q}^{j})$ in Algorithm 3 and Algorithm 4 is replaced by a set $\mathsf{U}({q}^{j})$ obtained by unirofmly sampling (with replacement) $|\mathsf{S}({q}^{j})|$ many elements from $\mathcal{L}({q}^{j})$, for all $q\in Q$ and for all $1\leq j\leq i$.
Lemma 6.0.

### 08 Apr 2024 v2

@@ 4.1. Correctness of Algorithm 1
Let $\mathsf{Bad}_{i}$ denote the event that the index $i$ is sampled (in Line 1) more that thresh number of times during the course of the algorithm, that is during the $t$ runs of the loop (Lines 1-1). And let the event $\mathsf{Bad}$ be $\cup_{i=1}^{k}\mathsf{Bad}_{i}$. We will now upper bound, $\Pr[\mathsf{Bad}_{i}]$. Let $W_{i}$ be the random variable counting the number times $i$ is sampled: $\mathbb{E}[W_{i}]=t\times\frac{\textsf{sz}_{i}}{\sum_{i=1}^{k}\textsf{sz}_{i}}$ which is less than $\frac{12(1+\epsilon_{\textsf{sz}})^{2}}{\epsilon^{2}}\log(4/\delta)$. By simple application of Chernoff Bound we see that
$\Pr[W_{i}\geq(1+\Delta)\mathbb{E}[W_{i}]]\leq exp\left(-\frac{\Delta^{2}\mathbb{E}[W_{i}]}{2+\Delta}\right)\leq exp\left(-\frac{\Delta\mathbb{E}[W_{i}]}{2}\right),$
@@
the last inequality is for $\Delta\geq 2$. For our purpose if we put $\Delta=\log(\frac{2k}{\delta})\frac{1}{\mathbb{E}[W_{i}]}$, since $|S_{i}|\geq\texttt{thresh}\geq\frac{12\cdot(1+\epsilon_{\textsf{sz}})^{2}}{\varepsilon^{2}}\cdot\log(\frac{4}{\delta})$ + $\log(\frac{2k}{\delta})$, we have $\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$. So by union bound $\Pr[\mathsf{Bad}]\leq\delta/2$. Observe that for large $k$, $\Delta\geq 2$ as desired. $\Pr[\mathsf{Bad_{i}}]\leq\frac{\delta}{2k}$.Observe that for large $k$, $\Delta\geq 2$ as desired. So by union bound, we have: ${\mathsf{Pr}}\!\left[\mathsf{Bad};{{\mathsf{U}}_{1,\ldots,k}}\right]=\Pr[\mathsf{Bad}]\leq\delta/2.$
4.2. Correctness of Algorithm 2
We prove the three parts of Theorem 2 individually.
Proof of Theorem 2(1)
@@ 4.2. Correctness of Algorithm 2
Consider an execution of Algorithm 2 and let $w=w_{1}w_{2}\ldots w_{\ell}\in\mathcal{L}({q}^{\ell})$ be the string constructed right before entering the branch at Line 2. Let $p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}$ be the value of the variable $\mathsf{pr_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}}}$ at level $i$ and let $v$ be the value of $\varphi$ before the function returns. Then we have,
@@
Thus, $v=\frac{\gamma_{0}}{\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}}{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}}=\gamma_{0}\cdot\prod\limits_{i=1}^{\ell}\frac{\textsf{sz}^{i}_{w_{i}}+\textsf{sz}^{i}_{1-w_{i}}}{\textsf{sz}^{i}_{w_{i}}}$
Recall that $P^{i}$ is the argument of the call at level $i$, and $P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}$ is the ${\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}$-predecessor of $P^{i}$. Observe that $\mathcal{L}(P^{i})=\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})\uplus\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})$, and thus, $|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}}})|+|\mathcal{L}(P^{i}_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}})|=|\mathcal{L}(P^{i})|$
@@
Now, observe that $|\mathcal{L}(P^{0})|=1$ and $P^{\ell}=\{{q}^{\ell}\}$.
Thus, under the event $\big{(}\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\big{)}=\texttt{Accurate}\texttt{AppUnion}$, we have
@@
It then follows that ${\mathsf{Pr}}\!\left[\mathtt{Fail}_{1}\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]$ is equal to 00.. Substituting, $v=\frac{\gamma_{0}}{p_{w_{\ell},\ell}\cdot p_{w_{\ell-1},\ell-1}\cdots p_{w_{1},1}}=\frac{\gamma_{0}}{\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}}$, we then have
${\mathsf{Pr}}\!\left[\mathtt{ret}_{\textnormal{{sample}}}=w\mid\texttt{Accurate}\texttt{AppUnion}\cap\texttt{AccurateN}_{\ell,q}\right]=v\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\frac{\gamma_{0}}{\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}}\cdot\prod_{i=\ell}^{1}p_{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},i}=\gamma_{0}$
Proof of Theorem 2(2)
@@
$=1{-}\!\!\!\!\!\!\sum\limits_{w\in\mathcal{L}({q}^{\ell})}\!\!\!\gamma_{0}=1-\gamma_{0}|\mathcal{L}({q}^{\ell})|\leq 1-\frac{2}{3e}\cdot\frac{|\mathcal{L}({q}^{\ell})|}{\mathsf{N}({q}^{\ell})}\leq 1-\frac{2}{3e}\cdot\frac{1}{(1+\beta)^{\ell}}\leq 1-\frac{2}{3e^{2}}$
Proof of Theorem 2(3)
@@
    $\texttt{thresh}\leq\frac{24\cdot e}{\beta^{2}}\log\left(\frac{16mn}{\eta}\right)\leq|\mathsf{S}({q}^{i-1})|.$
@@
    The event $\texttt{AccurateN}_{\leq\ell-1}$ demands that for each $i\leq\ell-1$, we have
    $\frac{|\mathcal{L}({q}^{i})|}{(1+\beta)^{i}}\leq\mathsf{N}({q}^{i})\leq(1+\beta)^{i}|\mathcal{L}({q}^{i})|.$
@@
$\begin{array}[]{l}{\mathsf{Pr}}\!\left[\bigcap_{\begin{subarray}{l}1\leq i\leq\ell\\ {\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}\in\{{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{0}},{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{1}}\}\end{subarray}}\texttt{Accurate}\texttt{AppUnion}_{i,{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@rgb@stroke{0.5}{0}{0}\pgfsys@color@rgb@fill{0.5}{0}{0}\mathtt{b}}}\;\;\Bigg{|}\;\ \texttt{AccurateN}_{\leq\ell-1};{{\mathsf{U}}_{\leq\ell-1}}\right]\geq 1-2n\cdot\eta^{\prime}=1-\frac{\eta}{2}\end{array}$
4.3. Correctness of Algorithm 3
We prove the main result (Theorem 3) by induction on the level $\ell$. For each level $\ell$, we first characterize the accuracy of the computed estimates $\{\mathsf{N}({q}^{\ell})\}_{q\in Q}$ and the quality of the sampled sets $\{\mathsf{S}({q}^{\ell})\}_{q\in Q}$, assuming that these were computed when the samples at the lower levels $11,2,\ldots,\ell-$ 1 are perfectly uniform and have size thresh. After this, we characterize the real computed estimates and samples by arguing that the distance of the corresponding random variables, from those computed using perfectly uniform samples, is small. We first establish some helpful auxiliary results (Lemma 1 and Lemma 2) towards our main proof. As before, we will often invoke distribution entanglement of ${\sf S}_{\leq\ell-1}$ by ${\mathsf{U}}_{\leq\ell-1}$ for each level $\ell$.
Lemma 1 ().